OpenStudy (anonymous):

Velocity vector of a fluid in a cylindrical vessel is of the form v=ω×r, where ω is the constant rotation vector. Show that div V = 0?

OpenStudy (anonymous):

Could you please clarify what you mean by "div V"?

OpenStudy (anonymous):

divergence of V

OpenStudy (anonymous):

Ah... further in Calc III than I know... sorry. :P

OpenStudy (anonymous):

not a problem. Thanks for looking

OpenStudy (nikvist):

\[\nabla\cdot\vec{v}=\nabla\cdot(\vec\omega\times\vec{r})=(\nabla\times\vec{\omega})\cdot\vec{r}-(\nabla\times\vec{r})\cdot\vec{\omega}\] \[\vec\omega-constant\quad\Rightarrow\quad\nabla\times\vec{\omega}=0\] \[\vec{r}-only\enspace radial\quad\Rightarrow\quad\nabla\times\vec{r}=0\] \[\Rightarrow\quad\nabla\cdot\vec{v}=0\]